\begin{thebibliography}{10}

\bibitem{AbdolrazaghiMMN}
F.~Abdolrazaghi and A.~Razani.
\newblock On the weak solutions of an overdetermined system of nonlinear
  fractional partial integro-differential equations.
\newblock {\em Miskolc Mathematical Notes}, 20(1):3--16, 2019.

\bibitem{agarwal2010survey}
R.~P. Agarwal, M.~Benchohra, and S.~Hamani.
\newblock A survey on existence results for boundary value problems of
  nonlinear fractional differential equations and inclusions.
\newblock {\em Acta Applicandae Mathematicae}, 109:973--1033, 2010.

\bibitem{aristov2017exact}
A.~I. Aristov.
\newblock On exact solutions of a sobolev equation.
\newblock {\em Mathematical Notes}, 101(5-6):928--941, 2017.

\bibitem{BehboudiFilmat}
F.~Bahboudi and A.~Razani.
\newblock Two weak solutions for a singular (p; q)- laplacian problem.
\newblock {\em Filomat}, Accepted.

\bibitem{barenblatt1960basic}
G.~Barenblatt, I.~P. Zheltov, and I.~Kochina.
\newblock Basic concepts in the theory of seepage of homogeneous liquids in
  fissured rocks [strata].
\newblock {\em Journal of Applied Mathematics and Mechanics}, 24(5):1286--1303,
  1960.

\bibitem{benchohra2009boundary}
M.~Benchohra, S.~Hamani, and S.~Ntouyas.
\newblock Boundary value problems for differential equations with fractional
  order and nonlocal conditions.
\newblock {\em Nonlinear Analysis: Theory, Methods \& Applications},
  71:2391--2396, 2009.

\bibitem{bonanno2011three}
G.~Bonanno and G.~M. Bisci.
\newblock Three weak solutions for elliptic dirichlet problems.
\newblock {\em Journal of Mathematical Analysis and Applications}, 382(1):1--8,
  2011.

\bibitem{bonanno2010structure}
G.~Bonanno and S.~A. Marano.
\newblock On the structure of the critical set of non-differentiable functions
  with a weak compactness condition.
\newblock {\em Applicable Analysis}, 89:1--10, 2010.

\bibitem{chu2017weak}
J.~Chu, S.~Heidarkhani, A.~Salari, and G.~Caristi.
\newblock Weak solutions and energy estimates for singular p-laplacian-type
  equations.
\newblock {\em Journal of Dynamical and Control Systems}, pages 1--13, 2017.

\bibitem{corvellec2010doubly}
J.-N. Corvellec, V.~Motreanu, and C.~Saccon.
\newblock Doubly resonant semilinear elliptic problems via nonsmooth critical
  point theory.
\newblock {\em Journal of Differential Equations}, 248:2064--2091, 2010.

\bibitem{davis1972quasilinear}
P.~L. Davis.
\newblock A quasilinear parabolic and a related third order problem.
\newblock {\em Journal of Mathematical Analysis and Applications},
  40(2):327--335, 1972.

\bibitem{dinmohammadi2017analytical}
A.~Dinmohammadi, A.~Razani, and E.~Shivanian.
\newblock Analytical solution to the nonlinear singular boundary value problem
  arising in biology.
\newblock {\em Boundary Value Problems}, 2017(1):63, 2017.

\bibitem{dinmohammadi2017existence}
A.~Dinmohammadi, E.~Shivanian, and A.~Razani.
\newblock Existence and uniqueness of solutions for a class of singular
  nonlinear two-point boundary value problems with sign-changing nonlinear
  terms.
\newblock {\em Numerical Functional Analysis and Optimization}, 38(3):344--359,
  2017.

\bibitem{EhsaniAA}
D.~Ehsani, M.~R. Mokhtarzadeh, and A.~Razani.
\newblock Expansions of dirichlet to neumann operators under perturbations.
\newblock {\em Applicable Analysis}, 89(5):789--799, 2010.

\bibitem{ewing1977coupled}
R.~E. Ewing.
\newblock A coupled non-linear hyperbolic-sobolev system.
\newblock {\em Annali di Matematica Pura ed Applicata}, 114(1):331--349, 1977.

\bibitem{GoodarziAAA}
Z.~Goodarzi and A.~Razani.
\newblock A periodic solution of the generalized forced li\'{e}nard equation.
\newblock {\em Abstract and Applied Analysis}, 2014(132450):5 pages, 2014.

\bibitem{goudarzi2019weak}
H.~Goudarzi, E.~Shivanian, and S.~H. Ghoncheh.
\newblock Weak solutions to a system of nonlinear fractional boundary value
  problems via variational form.
\newblock {\em Bulletin of the Malaysian Mathematical Sciences Society}, pages
  1--17, 2019.

\bibitem{heydari2018efficient}
M.~Heydari, H.~L. Dastjerdi, and M.~N. Ahmadabadi.
\newblock An efficient method for the numerical solution of a class of
  nonlinear fractional fredholm integro-differential equations.
\newblock {\em International Journal of Nonlinear Sciences and Numerical
  Simulation}, 19(2-3):165--173, 2018.

\bibitem{Khalkhali2012}
S.~M. Khalkhali, S.~Hidarkhani, and A.~Razani.
\newblock Infinitely many solutions for a fourth-order boundary-value problem.
\newblock {\em Electronic Journal of Differential Equations}, 164:1--14, 2012.

\bibitem{Khalkhali2013}
S.~M. Khalkhali and A.~Razani.
\newblock Multiple solutions for a quasilinear $(p,q)$-elliptic system.
\newblock {\em Electronic Journal of Differential Equations}, 2013(144):1--14,
  2013.

\bibitem{li2005existence}
F.~Li, Z.~Liang, and Q.~Zhang.
\newblock Existence of solutions to a class of nonlinear second order two-point
  boundary value problems.
\newblock {\em Journal of Mathematical Analysis and Applications},
  312:357--373, 2005.

\bibitem{MahdaviFilomat}
R.~Mahdavi~Khanghahi and A.~Razani.
\newblock Solutions for a singular elliptic problem involving the
  $p(x)$-laplacian.
\newblock {\em Filomat}, 32(14):4841--4850, 2018.

\bibitem{MakvandMJM}
M.~Makvand~Chaharlang, M.~A. Ragusa, and A.~Razani.
\newblock A sequence of radially symmetric weak solutions for some nonlocal
  elliptic problem in $\mathbb{R}^n$.
\newblock {\em Mediterranean Journal of Mathematics}, Accepted.

\bibitem{MakvandFilomat}
M.~Makvand~Chaharlang and A.~Razani.
\newblock Infinitely many solutions for a fourth order singular elliptic
  problem.
\newblock {\em Filomat}, 32(14):5003--5010, 2018.

\bibitem{MakvandCKMS}
M.~Makvand~Chaharlang and A.~Razani.
\newblock Existence of infinitely many solutions for a class of nonlocal
  problems with dirichlet boundary condition.
\newblock {\em Communications of the Korean Mathematical Society},
  34(1):155--167, 2019.

\bibitem{MakvandTJM}
M.~Makvand~Chaharlang and A.~Razani.
\newblock A fourth order singular elliptic problem involving $p$-biharmonic
  operator.
\newblock {\em Taiwanese Journal of Mathematics}, Accepted.

\bibitem{MakvandGMJ}
M.~Makvand~Chaharlang and A.~Razani.
\newblock Two weak solutions for some kirchhoff-type problem with neumann
  boundary condition.
\newblock {\em Georgian Mathematical Journal}, Accepted.

\bibitem{JeanMawhin}
J.~Mawhin.
\newblock {\em Critical Point Theory and Hamiltonian Systems}.
\newblock Springer-Verlag, Berlin/New York, 1989.

\bibitem{Mokhtarzadeh12}
M.~Mokhtarzadeh, M.~R. Pournaki, and A.~Razani.
\newblock An existence-uniqueness theorem for a class of boundary value
  problems.
\newblock {\em Fixed Point Theory}, 13(2):583--592, 2012.

\bibitem{nieto2009variational}
J.~J. Nieto and D.~O’Regan.
\newblock Variational approach to impulsive differential equations.
\newblock {\em Nonlinear Analysis: Real World Applications}, 10:680--690, 2009.

\bibitem{PournakiAML}
M.~R. Pournaki and A.~Razani.
\newblock On the existence of periodic solutions for a class of generalized
  forced li\'{e}nard equations.
\newblock {\em Applied Mathematics letters}, 20:248--254, 2007.

\bibitem{rabinowitz1986minimax}
P.~H. Rabinowitz.
\newblock {\em Minimax methods in critical point theory with applications to
  differential equations}.
\newblock American Mathematical Soc., 1986.

\bibitem{ragusa2004cauchy}
M.~A. Ragusa.
\newblock Cauchy--\textsc{D}irichlet problem associated to divergence form
  parabolic equations.
\newblock {\em Communications in Contemporary Mathematics}, 6(03):377--393,
  2004.

\bibitem{Razani2014}
A.~Razani.
\newblock An existence theorem for ordinary differential equation in menger
  probabilistic metric space.
\newblock {\em Miskolc Mathematical Notes}, 15(2):711--716, 2014.

\bibitem{shi1990initial}
D.~Shi.
\newblock On the initial boundary value problem of nonlinear the equation of
  the migration of the moisture in soil.
\newblock {\em Acta Mathematicae Applicatae Sinica}, 13(1):31--38, 1990.

\bibitem{tang2010some}
C.-L. Tang and X.-P. Wu.
\newblock Some critical point theorems and their applications to periodic
  solution for second order hamiltonian systems.
\newblock {\em Journal of Differential Equations}, 248:660--692, 2010.

\bibitem{ting1974cooling}
T.~W. Ting.
\newblock A cooling process according to two-temperature theory of heat
  conduction.
\newblock {\em Journal of Mathematical Analysis and Applications},
  45(1):23--31, 1974.

\bibitem{zhang2010positive}
S.~Zhang.
\newblock Positive solutions to singular boundary value problem for nonlinear
  fractional differential equation.
\newblock {\em Computers \& Mathematics with Applications}, 59:1300--1309,
  2010.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath,amssymb,amsthm}
\usepackage{cite}
\usepackage{enumitem}
\usepackage{authblk}
\usepackage{blindtext}
\usepackage[pagebackref=false,colorlinks,linkcolor=blue,citecolor=magenta]{hyperref}
\numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}{Definition}[section]
\newtheorem{remark}{Remark}[section]
\begin{document}
%\begin{frontmatter}
\title{Multiple weak solutions for a kind of time-dependent equation involving singularity}

\author{F. Abdolrazaghi\thanks{f.abdolrazaghi@edu.ikiu.ac.ir}}%fatemeh\_abdolrazaghi@yahoo.com

%\thanks{The first author was supported in part by the XXX Fund, Grant No.~YYYYY.}
\author{A. Razani\thanks{Corresponding author}\thanks{razani@sci.ikiu.ac.ir}}
\author{R. Mirzaei\thanks{r.mirzaei@sci.ikiu.ac.ir}}
\affil{Imam Khomeini International University\\ Faculty of  Science, Department of Pure Mathematics\\ Postal code: 34149-16818,
  Qazvin, Iran}

%\corresponding{author}
%\cortext[cor1]{Corresponding author}
\date{\empty}
\maketitle
\begin{abstract}
The existence of at least three weak solutions for a kind of nonlinear time-dependent
equation is studied. In fact, we consider the case that the source function has singularity at origin.
To this aim, the variational methods and the well-known critical points theorem are main tools.
\end{abstract}

\textbf{2010 Mathematics Subject Classification:}{35J20,\ 34B15}

\textbf{Keywords}: Sobolev equation, Weak solution, Critical point theory, Variational method, Singularity.

%\end{frontmatter}

\section{Introduction}
The linear Sobolev equations have a real physical background (\cite{barenblatt1960basic,shi1990initial,ting1974cooling}) and are studied in \cite{davis1972quasilinear,ewing1977coupled}. Because of their complexity,
%of the domain and the source functions
 they haven't exact solutions (except some very especial cases \cite{aristov2017exact}). There are different methods to study the solution of these problems. One of the standard methods is the fixed point theory that investigate the existence of solutions of nonlinear boundary
value problems \cite{agarwal2010survey,benchohra2009boundary,zhang2010positive,Mokhtarzadeh12,EhsaniAA,
PournakiAML,GoodarziAAA,Razani2014,dinmohammadi2017analytical,dinmohammadi2017existence}. The calculus of variation is another impressive technique and for using this technique, one needs to show that the given boundary value problem should possess a variational structure on some convenient spaces
\cite{AbdolrazaghiMMN,BehboudiFilmat,chu2017weak,corvellec2010doubly, heydari2018efficient,JeanMawhin,
Khalkhali2013,Khalkhali2012,MakvandFilomat, MahdaviFilomat, MakvandMJM, MakvandCKMS, MakvandTJM,MakvandGMJ,li2005existence,nieto2009variational,rabinowitz1986minimax,ragusa2004cauchy,tang2010some,goudarzi2019weak}.

In the present paper, we study the weak solutions of
\begin{equation}\label{In1}
\left\{
\begin{array}{l}
\frac{\partial u}{\partial t}-\frac{\partial(\triangle u)}{\partial t}=\mu f(x,t,u) \ \  in\  \Omega,\\ %\tage{In1}\\
u=0\ \ on\ \partial\Omega,\\% \tage{In2}\\
u(x,0)=g(x)  \ \ x\in\Omega,\\%  \tage{In3}
\end{array}
\right.
\end{equation}
where
$\Omega$ is a non-empty bounded open subset of $\mathbb{R}^N$
with  $\partial \Omega\in C^1$, $\mu$ is a positive parameter,
$f:\Omega\times\mathbb{R}^{+}\times\mathbb{R}\rightarrow\mathbb{R} $ is a Carath\'{e}odory function and
has a singularity at the origin with respect to the time variable and $g:\Omega\rightarrow \mathbb{R}$ vanishes on $\partial \Omega$.\\
The aim of this paper is to find an interval for $\mu$ for which the problem
\eqref{In1} admits at least three distinct weak solutions.

By integrating the first equation of \eqref{In1} we get
\begin{equation}\label{Co1}
\int_{0}^{t}\frac{\partial u(x,s)}{\partial s}ds-\int_{0}^{t}\frac{\partial\Delta u(x,s)}{\partial s}ds=\int_{0}^{t}\mu f(x,s,u)ds,
\end{equation}
or
\begin{equation}\label{Co001}
-\Delta u(x,t)=\mu F(x,t,u)-u(x,t)+g(x)-\Delta g(x),
\end{equation}
where
\begin{equation}\label{Co3}
F(x,t,u)=\int_{0}^{t}f(x,s,u)ds.
\end{equation}
The equation \eqref{Co001} is a time-dependent elliptic equation.

\begin{definition}
A function $u:\Omega\rightarrow \mathbb{R}$
is called a weak solution of
the problem
\eqref{In1}
if
$u\in H_{0}^{1}$
and
\begin{equation}\label{p2}
\begin{array}{rl}
\int_{\Omega}\nabla u(x,t)\cdot\nabla v(x)dx&-\mu\int_{\Omega}F(x,t,u(x))v(x)dx+\int_{\Omega}u(x,t)v(x)dx\\
&-\int_{\Omega}g(x)v(x)dx+\int_{\Omega}\Delta g(x)v(x)dx=0,
\end{array}
\end{equation}
for all
$v\in H_{0}^{1}$ and $t\geq0$.
\end{definition}
\begin{definition}\label{functional}
Define the functionals
$\varphi,\vartheta:H_{0}^{1}\rightarrow \mathbb{R}$
 by $\varphi(u):=\frac{1}{2}{\|u\|}^{2}$
and
\[
\begin{array}{rl}
\vartheta(u):=&\int_{\Omega}\widetilde{F}(x,t,u)dx-\frac{1}{2\mu}\int_{\Omega}\left(u(x,t)
\right)^2dx+\frac{1}{\mu}\int_{\Omega}g(x)u(x,t)dx\\ & \\
& \quad \quad -\frac{1}{\mu}\int_{\Omega}\Delta g(x)u(x,t)dx,
\end{array}
\]
respective, where $\widetilde{F}(x,t,\eta):=\int_{0}^{\eta}F(x,t,s)ds$.
\end{definition}
Notice that $\varphi$ and $\vartheta$ are well-defined and $ C^1$, $\varphi^\prime,\vartheta^\prime\in X^{*}$,
$\varphi^\prime(u)(v)=\int_{\Omega}\nabla u(x)\cdot\nabla v(x)dx$
and
\[
\begin{array}{rl}
\vartheta^\prime(u)(v)=&\int_{\Omega}F(x,t,u(x))v(x)dx-\frac{1}{\mu}
\int_{\Omega}u(x,t)v(x)dx\\
& \\
&\quad \quad +\frac{1}{\mu}\int_{\Omega}g(x)v(x)dx-\frac{1}{\mu}\int_{\Omega}\Delta g(x)v(x)dx.
\end{array}
\]
\begin{remark}
A critical point of  $I_\mu:=\varphi-\mu\vartheta$ is exactly a weak solution of
\eqref{In1}.
\end{remark}
%The Sobolev Embedding Theorem
%\cite{bonanno2011three}, says $H_0^1(\Omega)\hookrightarrow L^{2^*}(\Omega)$
% i.e. there exists
%$c\in \mathbb{R}^+$ such that for all $u \in H_0^1(\Omega)$
%\begin{equation}\label{p3}
%\|u\|_{ L^{2^*}(\Omega)}\leq c\|u\|,
%\end{equation}
%where
%\begin{equation}\label{p4}
%c:=\frac{1}{\sqrt{N(N-2)\pi}}\left(\frac{N!}{2\Gamma(N/2+1)}\right)^{\frac{1}{N}},
%\end{equation}
%$\Gamma(z)$ is the Gamma function and $2^*=2N/(N-2)$.
Fix $q\in [1,2^*[$, Embedding Theorem \cite{bonanno2011three} shows
$H_0^1(\Omega)\overset{c}\hookrightarrow L^{q}(\Omega)$, i.e.
there exists $c_q>0$ such that for all $u \in H_0^1(\Omega)$
\begin{equation}\label{p6}
\|u\|_{ L^{q}(\Omega)}\leq c_q\|u\|,
\end{equation}
where
\begin{equation}\label{p7}
c_q\leq\frac{meas(\Omega)^{\frac{2^*-q}{2^*q}}}{\sqrt{N(N-2)\pi}}\left(\frac{N!}{2\Gamma(N/2+1)}\right)^{\frac{1}{N}},
\end{equation}
$\Gamma$ is the Gamma function, $2^*=2N/(N-2)$
and $meas(\Omega)$ denotes the Lebesgue measure of $\Omega$.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Three weak solutions}
In this section the existence of at least three weak solutions for the problem \eqref{In1} is proved. Due to do this, we apply \cite[Theorem 3.6]{bonanno2010structure} which is given below
\begin{theorem}\label{p1}(see \cite{bonanno2010structure}, Theorem 3.6).
let
$X$
be a reflexive real Banach space,
$\Phi:X\rightarrow \mathbb{R}$
be a coercive, continuously Gateaux differentiable and sequentially weakly lower semicontinuous functional whose Gateaux derivative admits a continuous inverse on
$X^{*}$,
$\Psi:X\rightarrow \mathbb{R}$
be a continuously Gateaux differentiable functional whose Gateaux derivative is compact such that
$\Phi(0)=\Psi(0)=0.$
Assume that there exist
$r>0$
and
$\overline{x}\in X$,
with
$r<\Phi(\overline{x})$,
such that:
\begin{enumerate}
\item
$\frac{\sup_{\Phi(x)\leq r}\Psi(x)}{r}<\frac{\Psi(\overline{x})}{\Phi(\overline{x})};$
\item
for each
$\lambda\in\Lambda_{r}:=]\frac{\Phi(\overline{x})}{\Psi(\overline{x})},\frac{r}{\sup_{\Phi(x)\leq r}\Psi(x)}[$
the functional
$\Phi-\lambda\Psi$
is coercive.
\end{enumerate}
Then, for each
$\lambda \in \Lambda_{r}$,
the functional
$\Phi-\lambda\Psi$
has at least three distinct critical points in
$X$.
\end{theorem}
Set
\begin{equation}\label{p8}
\begin{array}{ll}
D:=\underset{x\in\Omega}\sup \ dist(x,\partial\Omega),&
\kappa:=\frac{D\sqrt{2}}{2{\pi}^{N/4}}\left(\frac{\Gamma(N/2+1)}{D^N-(D/2)^N}\right)^{\frac{1}{2}},\\
K_1:=\frac{2\sqrt{2}c_1(2^N-1)}{D^2},&
K_2:=\frac{2^{\frac{q+2}{2}}c_q^q(2^N-1)}{qD^2}.
\end{array}
\end{equation}
Now, we can state the main result.

\begin{theorem}\label{Maint}
Let
$f:\Omega\times\mathbb{R}^{+}\times\mathbb{R}\rightarrow\mathbb{R}$
be a Carath\'{e}odory function and $g:\Omega\rightarrow \mathbb{R}$ vanishes on $\partial \Omega$. Assume
\begin{itemize}
\item[(1)] There exist non-negative constants $m_1$,$m_2$ and $q \in ]1,\frac{2N}{N-2}[$ such that
$$F(x,t,s)\leq m_1+m_2\mid s\mid^{q-1}+\frac{1}{\mu}\left(s-g(x)+\Delta g(x)\right)$$
for all $(x,t,s)\in\Omega\times\mathbb{R}^{+}\times\mathbb{R}$.
\item[(2)]
$\widetilde{F}(x,t,\eta)\geq \frac{1}{\mu}\left(\frac{1}{2}\eta^2-\eta g(x)+\eta\Delta g(x)\right)$
for every $(x,t,\eta) \in \Omega\times\mathbb{R}^{+}\times\mathbb{R}$.
\item[(3)]
There exist positive constants $a$ and $b<2$ such that
$$\widetilde{F}(x,t,\eta)\leq a(1+|\eta|^{b})+\frac{1}{\mu}\left(\frac{1}{2}\eta^2-\eta g(x)+\eta\Delta g(x)\right).$$
\item[(4)]
There exist positive constants $\alpha$, $\beta$ with $\beta>\alpha\kappa$ such that
$$\frac{\inf_{x \in \Omega}\left(\widetilde{F}(x,t,\beta)-\frac{1}{\mu}\left(\frac{1}{2}\beta^2-\beta g(x)+\beta\Delta g(x)\right)\right)}{{\beta}^{2}}>m_1\frac{K_1}{\alpha}+m_2K_2{\alpha}^{q-2},$$
where $\kappa,K_1,K_2$ are given by \eqref{p8}.
\end{itemize}
Then the problem \eqref{In1} has at least three weak solutions in $H_{0}^{1}(\Omega)$, for each parameter $\mu$  belonging to
$\Lambda(\alpha,\beta):= \frac{2(2^N-1)}{D^2}\times \left(\delta_1,\delta_2\right)$,
where\\ $\delta_1:= \frac{{\beta}^{2}}{\inf_{x \in \Omega}\left(\widetilde{F}(x,t,\beta)-\frac{1}{\mu}\left(\frac{1}{2}\beta^2-\beta g(x)+\beta\Delta g(x)\right)\right)}$ and $\delta_2:=\frac{1}{m_1\frac{K_1}{\alpha}+m_2K_2\alpha^{q-2}}$.
\end{theorem}
\begin{proof}
Set $X:=H_{0}^{1}(\Omega)$ and define the functionals $\varphi(u)$ and
$\vartheta(u)$ by Definition \ref{functional}. Clearly, $\vartheta$ and $\varphi$ satisfy the assumptions of \cite[Theorem 3.6]{bonanno2010structure}.
By (1)
\begin{equation}\label{m1}
\widetilde{F}(x,t,\eta)\leq \frac{1}{\mu}\left(\frac{1}{2}\eta^{2}-\eta g(x)+\eta\Delta g(x)\right)+m_1|\eta|+m_2\frac{|\eta|^{q}}{q}
\end{equation}
for every
$(x,t,\eta)\in\Omega\times\mathbb{R}^{+}\times\mathbb{R}$. Thus
\[
\begin{array}{rl}
\vartheta(u):=& \int_{\Omega}\widetilde{F}(x,t,u)dx-\frac{1}{2\mu}\int_{\Omega}\left(u(x,t)\right)^2dx+\frac{1}{\mu}\int_{\Omega}g(x)u(x,t)dx\\& \\
&\quad -\frac{1}{\mu}\int_{\Omega}\Delta g(x)u(x,t)dx\\ & \\
\leq& \frac{1}{\mu}\int_{\Omega}\left(\frac{1}{2}(u(x,t))^{2}-u(x,t) g(x)+u(x,t)\Delta g(x)\right)dx \\ & \\
& \quad +\int_{\Omega}\left(m_1|u(x,t)|+m_2\frac{|u(x,t)|^{q}}{q}\right)dx-\frac{1}{2\mu}\int_{\Omega}\left(u(x,t)\right)^2dx
\\ & \\
&\quad +\frac{1}{\mu}\int_{\Omega}g(x)u(x,t)dx-\frac{1}{\mu}\int_{\Omega}\Delta g(x)u(x,t)dx\\ & \\
\leq& m_1\parallel u\parallel_{L^1(\Omega)}+\frac{m_2}{q}\parallel u\parallel_{L^q(\Omega)}^q.
\end{array}
\]
Let
$r \in ]0,+\infty[$ such that $ \varphi(u)\leq r$. By \eqref{p6},
\[
\vartheta(u)\leq\left(\sqrt{2r}c_1m_1+\frac{2^{\frac{q}{2}}c_q^qm_2}{q}r^{\frac{q}{2}}\right).
\]
Set $\chi(r):=\frac{\sup_{u\in\varphi^{-1}]-\infty,r[}\vartheta(u)}{r}$.
Consequently
\begin{equation}\label{m3}
\chi(r)\leq\left(\sqrt{\frac{2}{r}}c_1m_1+\frac{2^{\frac{q}{2}}c_q^qm_2}{q}r^{\frac{q}{2}-1}\right),
\end{equation}
for every $r>0$.

By \eqref{p8}, there is $x_0\in\Omega$ such that $B(x_0,D)\subseteq\Omega$. Set
\begin{eqnarray}
u_\beta(x,t):=\left\{
  \begin{array}{ll}
  0 \ \ \ \ \  x \in\Omega\backslash B(x_0,D), \\
  \frac{2\beta}{D}(D-|x-x_0|) \ \ x\in\ B(x_0,D)\backslash B(x_0,D/2),\\
\beta  \ \ \ \ \  x\in B(x_0,D/2).
  \end{array}
\right.
\end{eqnarray}
Thus
$u_{\beta}\in H_0^1(\Omega)$.
So
\begin{eqnarray}\label{m5}
\nonumber \varphi(u_\beta)&=&\frac{1}{2}\int_{\Omega}|\nabla u_\beta(x,t) |^2dx \\
\nonumber&=&\frac{1}{2}\int_{B(x_0,D)\backslash B(x_0,D/2)}\frac{(2\beta)^2}{D^2}dx \\
\nonumber&=&\frac{1}{2}\frac{(2\beta)^2}{D^2}(meas(B(x_0,D))-meas(B(x_0,D/2)))\\
&=&\frac{1}{2}\frac{(2\beta)^2}{D^2}\frac{\pi^{N/2}}{\Gamma(N/2+1)}\left(D^N-(D/2)^N\right).
\end{eqnarray}
If we force
$\beta>\alpha\kappa$, by (4), $\alpha^2<\varphi(u_{\beta})$ because $\alpha^2<\frac{\beta^2}{\kappa^2}$. Also by assumption (2),
\begin{equation}\label{m6}
\begin{array}{rl}
\vartheta(u_\beta):=&\int_{\Omega}\widetilde{F}(x,t,u_\beta)dx-
\frac{1}{2\mu}\int_{\Omega}\left(u_\beta(x,t)\right)^2dx+\frac{1}{\mu}\int_{\Omega}g(x)u_\beta(x,t)dx\\ & \\
&\quad \quad -\frac{1}{\mu}\int_{\Omega}\Delta g(x)u_\beta(x,t)dx\\ & \\ =&\int_{\Omega}\left[\widetilde{F}(x,t,u_\beta)-\frac{1}{\mu}\left(\frac{1}{2}u_\beta(x,t)^2-g(x)u_\beta(x,t)+\Delta g(x)u_\beta(x,t)\right)\right]dx\\ & \\
\geq&\int_{B(x_0,D/2)}\left[\widetilde{F}(x,t,u_\beta)-\frac{1}{\mu}\left(\frac{1}{2}u_\beta(x,t)^2-g(x)u_\beta(x,t)+\Delta g(x)u_\beta(x,t)\right)\right]dx\\ & \\
\geq &\inf_{x\in\Omega}\left(\widetilde{F}(x,t,\beta)-\frac{1}{\mu}\left(\frac{1}{2}\beta^2-\beta g(x)+\beta\Delta g(x)\right)\right)\frac{\pi^{N/2}}{\Gamma(N/2+1)}\frac{D^N}{2^N}.
\end{array}
\end{equation}
Next by dividing \eqref{m5} on \eqref{m6}, we have
\begin{equation}\label{m7}
\frac{\vartheta(u_\beta)}{\varphi(u_\beta)}\geq \frac{D^2}{2(2^N-1)}\frac{\inf_{x\in\Omega}\left(\widetilde{F}(x,t,\beta)-\frac{1}{\mu}\left(\frac{1}{2}\beta^2-\beta g(x)+\beta\Delta g(x)\right)\right)}{\beta^2}.
\end{equation}
Using \eqref{m3}, assumption (4) implies
\begin{eqnarray}\label{m8}
\nonumber\chi(\alpha^2)&\leq&(\frac{\sqrt{2}c_1m_1}{\alpha}+\frac{2^{\frac{q}{2}}c_q^q m_2\alpha^{q-2}}{q})\\
\nonumber&=&\frac{D^2}{2(2^N-1)}(m_1\frac{K_1}{\alpha}+m_2K_2{\alpha}^{q-2})\\
\nonumber&<&\frac{D^2}{2(2^N-1)}\frac{\inf_{x\in\Omega}(\widetilde{F}(x,t,\beta)-U(x,t)-G(x)-\overset{\Delta}G(x))}{\beta^2}\\
\nonumber&\leq&\frac{\vartheta(u_\beta)}{\varphi(u_\beta)}.
\end{eqnarray}
Assuming $b<2$ and considering $|u|^b\in L^{\frac{2}{s}}(\Omega)$ for all $u \in X$,
 H\"{o}lder's inequality for $ u \in X$ implies
$\int_{\Omega}|u(x,t)|^b dx\leq\parallel u\parallel_{L^2(\Omega)}^b (meas(\Omega))^{\frac{2-b}{2}}$.
Therefore equation \eqref{p6} shows for all $u \in X$
\[
\int_{\Omega}|u(x,t)|^b dx\leq c_2^b\parallel u\parallel^b (meas(\Omega))^{\frac{2-b}{2}},
\]
and by assumption (3),
\[
\begin{array}{rl}
I_\mu(u)=& \varphi(u)-\mu\vartheta(u)\\ & \\
=& \frac{\parallel u\parallel^2}{2}-\mu\int_{\Omega}\widetilde{F}(x,t,u)dx+
\frac{1}{2}\int_{\Omega}\left(u(x,t)\right)^2dx\\&\\
&\quad -\int_{\Omega}g(x)u(x,t)dx+\int_{\Omega}\Delta g(x)u(x,t)dx\\&\\
\geq & \frac{\parallel u\parallel^2}{2}-\mu\int_{\Omega}a\left(1+|u(x,t)|^{b}\right)dx\\ &\\
\geq &\frac{\parallel u\parallel^2}{2}-\mu a c_2^b (meas(\Omega))^{\frac{2-b}{2}}\parallel u\parallel^b-a\mu meas(\Omega).
\end{array}
\]
This means for every $
\mu \in \Lambda(\alpha,\beta)\subseteq \left] \frac{\vartheta(u_\beta)}{\varphi(u_\beta)},\frac{\alpha^2}{\sup_{\varphi(u)\leq \alpha^2}\vartheta(u)}\right[$,
$I_\mu$ is coercive.
Therefore by Theorem \ref{p1} for each
$\mu \in \Lambda(\alpha,\beta)$
the functional $I_\mu$
has at least three distinct critical points that they are weak solutions of the problem
\eqref{In1}.
\end{proof}
%----------------------------------------------------------------------------------------------------------------------------------------------------------
\section{Numerical Experiment}
Now, we present an example.

\begin{equation}\label{N1}
\left\{
\begin{array}{l}
  \frac{\partial u}{\partial t}-\frac{\partial(\Delta u)}{\partial t}=\frac{1}{100}\frac{99}{100t}\left(1+\frac{\exp(-t)}{99}\right)(8+100u+u^2) \ \in \Omega , %\label{N1}\\
  u\mid_{\partial\Omega}=0,\\ %\label{N2}\\
  u(x,0)=\frac{1}{1000}\left(\frac{1}{100}-\left(x_1^2+x_2^2+x_3^2\right)\right) \ \ x\in\Omega,
\end{array}
\right.
\end{equation}
where $\Omega :=\left\{(x_1,x_2,x_3)\in \mathbb{R}^3, x_1^2+x_2^2+x_3^2\leq0.1\right\}$, then $\mu=0.01, N=3, D=r=0.1, 2^*=6,$ $g(x)=0.001\left(0.01-\left(x_1^2+x_2^2+x_3^2\right)\right),$ $\Delta g(x)=-0.006$ and
$f(x,t,u)=\frac{99}{100t}\left(1+\frac{\exp(-t)}{99}\right)(8+100u+u^2)$. Now, setting $q=3$, then
\[
\begin{array}{l}
c_1\leq0.00445759,\quad  c_q\leq0.171543,\\
\kappa=1.16798,\quad  K_1\leq8.82557,\quad  K_2\leq6.66307.
\end{array}
\]
Clearly $F(x,t,s)=\frac{99}{100}\left(1+\frac{\exp(-t)}{99}\right)\left(8+100s+s^2\right)$, suppose
$m_1=9$ and $m_2=1$, then the assumption (1) of the Theorem \ref{Maint} is satisfied, i.e.
\[
\begin{array}{l}
\frac{99}{100}\left(1+\frac{\exp(-t)}{99}\right)\left(8+100s+s^2\right)\leq \\
9+ s^{2}+\frac{1}{0.01}\left(s-0.001\left(0.01-\left(x_1^2+x_2^2+x_3^2\right)\right)-0.006\right),
\end{array}
\]
for all $(x,t,s)\in\Omega\times\mathbb{R}^{+}\times\mathbb{R}$.\\ Obviously $\widetilde{F}(x,t,\eta)=\frac{99}{100}\left(1+\frac{\exp(-t)}{99}\right)\left(8\eta+50\eta^2+\frac{\eta^3}{3}\right)$,
then it can be easily verified that the assumption (2) of the Theorem \ref{Maint} holds, i.e.
for all $(x,t,s)\in\Omega\times\mathbb{R}^{+}\times\mathbb{R}$
\[
\begin{array}{l}
\frac{99}{100}\left(1+\frac{\exp(-t)}{99}\right)
\left(8\eta+50\eta^2+\frac{\eta^3}{3}\right)\geq \\
\frac{1}{0.01}\left(\frac{1}{2}\eta^2- 0.001\eta\left(0.01-\left(x_1^2+x_2^2+x_3^2\right)\right)-0.006\eta\right).
\end{array}
\]
Also, by choosing $a=b=10$, the assumption (3) of the Theorem \ref{Maint} is satisfied, i.e.
for all $(x,t,s)\in\Omega\times\mathbb{R}^{+}\times\mathbb{R}$
\[
\begin{array}{l}
\frac{99}{100}\left(1+\frac{\exp(-t)}{99}\right)\left(8\eta+50\eta^2+\frac{\eta^3}{3}\right)\leq\\
10(1+\eta^{10})+\frac{1}{0.01}\left(\frac{1}{2}\eta^2- 0.001\eta\left(0.01-\left(x_1^2+x_2^2+x_3^2\right)\right)-0.006\eta\right).
\end{array}
\]
More, set $\alpha=1$ and $\beta=500>\alpha\kappa$ hence, for all $t\geq 0$, it is not difficult to see that
\begin{eqnarray}
\nonumber
162.872&=&\frac{\inf_{x \in \Omega}\left\{\left(
\begin{array}{c}
\frac{99}{100}\left(1+\frac{\exp(-t)}{99}\right)\left(8\eta+50\eta^2+\frac{\eta^3}{3}\right)- \\
\frac{1}{0.01}\left(\frac{1}{2}\eta^2- 0.001\eta\left(0.01-\left(x_1^2+x_2^2+x_3^2\right)\right)-0.006\eta\right) \\
\end{array}
\right)
\right\}}{{\beta}^{2}}\\ \nonumber&>&m_1K_1+m_2K_2=86.0932.
\end{eqnarray}
Furthermore, it is observed that
$\mu=0.01\in\left]\frac{1}{162.872}%=0.0060784,0.0116153=
,\frac{1}{86.0932}\right[$,
therefore the problem \eqref{N1} admits at least three week solutions in according to the Theorem \ref{Maint}.
%*********************************Bibliography******************************************************

\textbf{Acknowledgements}\\
The authors are very grateful to anonymous reviewers for carefully reading the paper and for their comments and
suggestions which have improved the paper very much.
%\section*{References}
\bibliographystyle{abbrv}
\bibliography{myref}
%-------------------------------------------------------------------------------------------------------
\end{document}